OpenStudy (anonymous):

Complete the following fractions so that it is close to but less than 1/2 ?/100

OpenStudy (anonymous):

figure out what the values of the ? should be to equal exactly 1/2 and then subtract one from that value. Think about money and half dollars.

OpenStudy (anonymous):

Please Tell me What it is what is closer but less than 1/2 but the first question is /100

sam (.sam.):

You know that \[\frac{ 50 }{ 100 }=\frac{ 1 }{ 2 }\] Substract 50 by 1 =49 \[\frac{49}{100}\]

OpenStudy (anonymous):

So Its 49/100?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{x}{100}<\frac{1}{2} \]\[x<50 \]